# SurgiATM: A Physics-Guided Plug-and-Play Model for Deep Learning-Based Smoke Removal in Laparoscopic Surgery

### v1

@@ 4.1 Implementation Details
By default, we set the smoothing factor $\eta=0.1$ and the window size $z=15$ [Kaiming_2011_Single_Image]. All experiments are conducted on a 24GB NVIDIA RTX 3090 GPU. To ensure fair comparisons, training configurations are largely standardized across all methods, employing the Adam optimizer with a learning rate of 0.0002 and epochs of 50. Due to GPU memory constraints, the batch size depends on their respective publications.
@@ 4.2.1 Datasets
In this study, we utilize three datasets: Cholec80 [Twinanda_2017_EndoNet_A], VASST-desmoke [Xia_2024_A_New, Xia_2025_In_Vivo], and the Hamlyn Centre Laparoscopic and Endoscopic Dataset Videos (referred to as the Hamlyn Dataset for simplicity) [Giannarou_2013_Probabilistic_Tracking, Ye_2017_Self_Supervised]. The Hamlyn Dataset is specifically reserved for external evaluation. For Cholec80, 500 smokeless and 500 smoky frames are randomly sampled from each video [Leibetseder_2017_Real_Time], with the smoke-free frames then blended with synthetic smoke [Bolkar_2018_Deep_Smoke] for training. VASST-desmoke is a recently released paired benchmark for laparoscopic surgical smoke removal. It provides both real smoky frames and corresponding ground-truth smoke-free annotations, thereby enabling evaluation on real surgical smoke with intuitive full-reference metrics. Potentially, VASST-desmoke also allows models to learn from real smoke. The Hamlyn Dataset is excluded from all training and statistical analysis procedures in this work, and its frames are sampled from the videos at 25 frames per second (FPS). We perform five-fold cross-evaluation, ensuring no video overlap between the training and evaluation sets. The three datasets record various surgical procedures with inconsistent camera devices and lighting environments in different hospitals, significantly challenging the stability and generalizability of surgical desmoking methods.
@@ 4.2.2 Metrics
For the evaluation of paired datasets, we leverage full-reference metrics: CIEDE2000, PSNR, RMSE, and SSIM, to evaluate perceptual color difference, reconstruction quality, prediction error, and structure similarity, respectively. For the unpaired datasets, we evaluate model performance using non-reference metrics, including BRISQUE and NIQE to measure how close an image is to natural scenes, and FADE [Choi_2015_Referenceless_Prediction], which specifically measures fog-related degeneration.
Table 2: Training and Evaluation on VASST-desmoke (Baseline / Baseline + SurgiATM)
Method Year Backbone Loss Function CIEDE2000 $\downarrow$ PSNR $\uparrow$ RMSE $\downarrow$ SSIM $\uparrow$
AODNet [Li_2017_AOD_Net] 2017 CNN MSE 10.349 / 6.327 17.378 / 21.994 0.147 / 0.090 0.693 / 0.789
GCANet [Chen_2019_Gated_Context] 2019 CNN MSE 7.893 / 5.369 21.091 / 23.907 0.095 / 0.073 0.729 / 0.806
De-smokeGCN [Chen_2020_De_smokeGCN] 2020 U-Net MAE + Others 9.577 / 7.348 19.779 / 22.127 0.115 / 0.094 0.550 / 0.717
MSBDN [Dong_2020_Multi_Scale] 2020 U-Net MSE 6.679 / 5.164 21.604 / 23.881 0.088 / 0.073 0.651 / 0.813
SVPNet [Wang_2024_Smoke_veil] 2024 U-Net MSE + Others 5.574 / 5.141 23.464 / 24.362 0.073 / 0.070 0.789 / 0.820
CGAN-DC [Chen_2019_Gated_Context] 2019 cGAN Adversarial + Others 6.735 / 5.812 22.453 / 23.809 0.083 / 0.079 0.676 / 0.782
MARS-GAN [Hong_2023_MARS_GAN] 2023 CycleGAN Adversarial + Others 6.389 / 5.899 20.121 / 23.488 0.113 / 0.076 0.602 / 0.819
SSIM-PAN [Sidorov_2020_Generative_Smoke] 2020 GAN Adversarial + Others 7.046 / 5.811 21.465 / 23.477 0.090 / 0.075 0.655 / 0.750
RSTN [Wang_2023_Surgical_smoke] 2023 SwinT Charbonnier + Others 5.385 / 4.976 23.263 / 24.306 0.074 / 0.071 0.750 / 0.824
LGUTransformer [Wang_2024_Desmoking_of] 2024 SwinT + U-Net Charbonnier 9.927 / 5.990 18.238 / 23.376 0.121 / 0.080 0.510 / 0.761
Table 3: Training on Cholec80 with synthetic smoke and evaluation on VASST-desmoke
@@
AODNet [Li_2017_AOD_Net] 9.609 / 7.774 18.621 / 21.526 0.132 / 0.109 0.729 / 0.798
GCANet [Chen_2019_Gated_Context] 9.803 / 7.042 19.338 / 22.220 0.119 / 0.097 0.747 / 0.801
De-smokeGCN [Chen_2020_De_smokeGCN] 8.370 / 7.585 20.654 / 22.088 0.109 / 0.102 0.716 / 0.788
MSBDN [Dong_2020_Multi_Scale] 8.277 / 7.327 20.360 / 22.153 0.114 / 0.103 0.769 / 0.810
SVPNet [Wang_2024_Smoke_veil] 8.105 / 7.797 21.196 / 21.712 0.110 / 0.109 0.806 / 0.793
CGAN-DC [Chen_2019_Gated_Context] 8.177 / 7.525 20.663 / 21.988 0.113 / 0.105 0.770 / 0.805
MARS-GAN [Hong_2023_MARS_GAN] 7.808 / 7.444 21.444 / 22.110 0.107 / 0.104 0.792 / 0.811
SSIM-PAN [Sidorov_2020_Generative_Smoke] 8.350 / 7.600 20.338 / 21.805 0.115 / 0.106 0.761 / 0.806
RSTN [Wang_2023_Surgical_smoke] 7.963 / 7.533 21.034 / 22.065 0.111 / 0.107 0.763 / 0.813
LGUTransformer [Wang_2024_Desmoking_of] 7.357 / 7.352 22.108 / 22.111 0.105 / 0.104 0.812 / 0.818
Table 4: Training with synthetic smoke and evaluation with real smoke on Cholec80
@@
AODNet [Li_2017_AOD_Net] 12.551 / 14.814 0.440 / 0.484 10.449 / 9.434
GCANet [Chen_2019_Gated_Context] 7.795 / 11.566 0.407 / 0.403 9.866 / 9.331
De-smokeGCN [Chen_2020_De_smokeGCN] 4.857 / 5.575 0.384 / 0.376 10.885 / 9.569
MSBDN [Dong_2020_Multi_Scale] 11.838 / 12.176 0.421 / 0.411 10.299 / 9.026
SVPNet [Wang_2024_Smoke_veil] 12.812 / 12.134 0.416 / 0.402 10.579 / 9.387
CGAN-DC [Chen_2019_Gated_Context] 10.258 / 9.945 0.413 / 0.404 9.517 / 9.039
MARS-GAN [Hong_2023_MARS_GAN] 13.999 / 11.680 0.416 / 0.413 13.336 / 13.884
SSIM-PAN [Sidorov_2020_Generative_Smoke] 10.539 / 11.093 0.411 / 0.403 9.811 / 11.999
RSTN [Wang_2023_Surgical_smoke] 13.868 / 12.652 0.420 / 0.414 12.245 / 16.910
LGUTransformer [Wang_2024_Desmoking_of] 9.126 / 11.236 0.417 / 0.416 13.734 / 12.471
Table 5: Training on VASST-desmoke and evaluation on Hamlyn Dataset
@@
AODNet [Li_2017_AOD_Net] 36.549 / 30.515 0.331 / 0.322 12.418 / 11.851
GCANet [Chen_2019_Gated_Context] 13.354 / 26.555 0.324 / 0.294 11.512 / 10.983
De-smokeGCN [Chen_2020_De_smokeGCN] 33.236 / 28.007 0.329 / 0.325 29.448 / 18.802
MSBDN [Dong_2020_Multi_Scale] 26.992 / 25.198 0.384 / 0.322 14.224 / 10.944
SVPNet [Wang_2024_Smoke_veil] 39.862 / 30.325 0.401 / 0.317 16.245 / 13.306
CGAN-DC [Chen_2019_Gated_Context] 14.052 / 16.699 0.352 / 0.336 16.616 / 13.216
MARS-GAN [Hong_2023_MARS_GAN] 33.722 / 30.601 0.377 / 0.369 15.011 / 16.381
SSIM-PAN [Sidorov_2020_Generative_Smoke] 11.936 / 16.047 0.357 / 0.298 16.613 / 20.297
RSTN [Wang_2023_Surgical_smoke] 35.629 / 30.232 0.412 / 0.330 16.442 / 12.645
LGUTransformer [Wang_2024_Desmoking_of] 35.260 / 29.706 0.295 / 0.294 15.300 / 12.135
Table 6: Training on Cholec80 with synthetic smoke and evaluation on Hamlyn Dataset
@@
AODNet [Li_2017_AOD_Net] 30.292 / 31.168 0.426 / 0.432 11.074 / 11.834
GCANet [Chen_2019_Gated_Context] 19.366 / 26.102 0.385 / 0.293 11.628 / 11.083
De-smokeGCN [Chen_2020_De_smokeGCN] 13.920 / 17.647 0.330 / 0.315 11.628 / 11.467
MSBDN [Dong_2020_Multi_Scale] 26.630 / 27.559 0.409 / 0.332 12.152 / 11.882
SVPNet [Wang_2024_Smoke_veil] 29.065 / 28.375 0.406 / 0.267 14.358 / 12.292
CGAN-DC [Chen_2019_Gated_Context] 25.175 / 24.127 0.408 / 0.315 10.445 / 10.741
MARS-GAN [Hong_2023_MARS_GAN] 29.811 / 27.287 0.409 / 0.311 16.443 / 15.136
SSIM-PAN [Sidorov_2020_Generative_Smoke] 24.605 / 26.001 0.395 / 0.312 10.852 / 19.337
RSTN [Wang_2023_Surgical_smoke] 26.305 / 28.347 0.414 / 0.328 13.875 / 14.217
LGUTransformer [Wang_2024_Desmoking_of] 23.974 / 26.497 0.383 / 0.346 16.786 / 17.448
@@ 4.3 Comparison of Quantitative Results
The quantitative results on VASST-desmoke are presented in Tables 2 and 3. They demonstrate that incorporating SurgiATM effectively improves model performance, regardless of the backbone architecture or loss function. However, for SVPNet in Table 3, SurgiATM yields no improvement in SSIM. We argue that, compared to other desmoking methods, SVPNet [Wang_2024_Smoke_veil] employs an energy function [Wang_2018_Variational_based] for restoration. This is a physics-guided approach similar to ours, and its three physics-related parameters are predicted by three independent U-Nets. As a result, incorporating an extra physics-guided component (our SurgiATM) into SVPNet may slightly disrupt the internal physical consistency, leading to the observed deterioration in SSIM.
For Cholec80, we report the evaluation results in Table 4. Our SurgiATM effectively improves most model performances in terms of the fog-related metric (i.e., FADE score). In terms of the external evaluation on the Hamlyn Dataset, as shown in Tables 6 and 5, for the defogging metric FADE, employing our SurgiATM enhances most methods, except for the traditional dehazing method AODNet in Table 6. We attribute this to two factors: 1) the simple architecture and insufficient trainable weights of AODNet (i.e., only five convolutional layers) compared to other methods, and 2) the non-Laplacian error distribution of AODNet when trained on Cholec80 with synthetic smoke, as reported in Table 1. In contrast, although GCANet and SSIM-PAN also exhibit non-Laplacian error distributions in Cholec80, and CGAN-DC and MARS-GAN in VASST-desmoke, their performances are effectively enhanced owing to both their well-designed architectures and the strong generalizability of SurgiATM.
On the other hand, SurgiATM demonstrates limited effectiveness in terms of BRISQUE and NIQE across the ten existing methods in Tables 4, 5, and 6. The two possible reasons are: 1) these two metrics are designed to measure similarity to natural images, which may not generalize well to surgical endoscopic frames due to the substantial domain gap between them [Wang_2024_Desmoking_of]; and 2) the Laplacian-guided error minimization of SurgiATM may contribute to dissimilarity from natural images (Gaussian-like error).
Figure 3: Error comparison in VASST-desmoke. “+” indicates the baseline integrated with our SurgiATM. ”GT” represents ground truth. The methods are trained in Cholec80 with synthetic smoke. Orange boxes mark the errors declined by SurgiATM. Overall, “Baseline + SurgiATM” exhibits more accurate restoration.
Figure 4: Comparison of SurgiATM with baselines on real datasets: Cholec80 (left two columns) and the Hamlyn Dataset (right two columns). Blue boxes highlight the comparatively accurate and stable restoration achieved by SurgiATM.
@@ 4.4 Comparison of Qualitative Results
In this section, we present qualitative comparisons across several representative baselines: De-smokeGCN, GCANet, MARS-GAN, and SVPNet, each selected for different reasons. Specifically, De-smokeGCN [Chen_2020_De_smokeGCN] is a classic DNN-based desmoking method introduced in 2020; GCANet [Chen_2019_Gated_Context] is designed for natural dehazing and deraining; MARS-GAN represents an adversarial-learning-based approach; and SVPNet [Wang_2024_Smoke_veil], published in 2024, uses a physics-based loss function tailored for surgical desmoking. We train the models on Cholec80 with synthetic smoke and evaluate them on VASST-desmoke, Cholec80, and the Hamlyn Dataset with real surgical smoke, challenging their stability and generalizability.
As shown in the error maps in Fig. 3, SurgiATM effectively reduces desmoking errors in a computationally efficient and interpretable manner, making it well-suited for deployment on edge devices and for stable surgical desmoking. For the natural dehazing method GCANet, it causes color distoration in Fig. 3(a), which is corrected to a certain degree by using SurgiATM. In addition, the proposed approach can mitigate the grid artifact (see the first two rows in Fig. 3), demonstrating promising stability. As shown in Fig. 4, the desmoking results of the baselines are improved to varying degrees. For instance, De-smokeGCN and GCANet can produce more stable results with SurgiATM by reducing the grid artifacts (see Fig. 4(a)-(d) of De-smokeGCN) and correcting the restoration colors (see Fig. 4(d) of GCANet). Moreover, SurgiATM helps baselines eliminate more smoke, as shown in Fig. 4(a) for MARS-GAN and Fig. 4(c) for SVPNet, GCANet, and MARS-GAN.

### v2

@@ 4.1 Implementation Details
For all comparisons, we set the smoothing factor $\eta=0.1$ because this value contributes to better results in out ablation study; and set the window size $z=15$ inherited from DCP (Kaiming et al., 2011). The frames are resized into $\times256$. All experiments are conducted on a 24GB NVIDIA RTX 3090 GPU. SurgiATM is extended as a output layer at the end of all approaches, without any modification to original architectures, input forms, loss functions and hyper-parameters. To ensure fair comparisons, training configurations are standardized across all methods, employing the Adam optimizer with a learning rate of 0.0002, beta1 of 0.9, beta2 of 0.999, and epochs of 50. Due to GPU memory constraints, the batch size depends on their respective publications. Please visit to https://github.com/MingyuShengSMY/SurgiATM for more implementation details. We implement SurgiATM as a torch.nn.Module and upload it onto the link, which can be directly used.
@@ 4.2.1 Datasets
In this study, we utilize three datasets: Cholec80 (Twinanda et al., 2017), VASST-desmoke (Xia et al., 2024, 2025), and the Hamlyn Centre Laparoscopic and Endoscopic Dataset Videos (referred to as the Hamlyn Dataset for simplicity) (Giannarou et al., 2013; Ye et al., 2017). The Hamlyn Dataset is specifically reserved for external evaluation. For Cholec80, 500 smokeless and 500 smoky frames are randomly sampled from each video (Leibetseder et al., 2017), with the smoke-free frames then blended with synthetic smoke (Bolkar et al., 2018) for training. VASST-desmoke is a recently released paired benchmark for laparoscopic surgical smoke removal. It provides both real smoky frames and corresponding ground-truth smoke-free annotations, thereby enabling evaluation on real surgical smoke with intuitive full-reference metrics. Potentially, VASST-desmoke also allows models to learn from real smoke. In addition, Xia et al. (2025) further expand the VASST-desmoke, producing more paired frames sourced from Cholec80. For convenience, we note the latest part of the dataset as ”VASST-desmoke-C”. The Hamlyn Dataset is excluded from all training and statistical analysis procedures in this work, and its frames are sampled from the videos at 25 frames per second (FPS). We perform five-fold cross-evaluation, ensuring no video overlap between the training and evaluation sets. The three datasets record various surgical procedures with inconsistent camera devices and lighting environments in different hospitals, significantly challenging the stability and generalizability of surgical desmoking methods.
@@ 4.2.2 Metrics
For the evaluation of paired datasets, we leverage full-reference metrics: CIEDE2000, PSNR, RMSE, and SSIM, to evaluate perceptual color difference, reconstruction quality, prediction error, and structure similarity, respectively. For the unpaired datasets, we evaluate model performance using non-reference metrics, including BRISQUE and NIQE to measure how close an image is to natural scenes, and FADE (Choi et al., 2015), which specifically measures fog-related degeneration.
Table 5: Training and evaluation on VASST-desmoke-C (real smoke).
@@
AODNet (Li et al., 2017) 7.344 / 5.297 20.727 / 24.655 0.098 / 0.065 0.860 / 0.900
GCANet (Chen et al., 2019) 7.453 / 4.291 22.566 / 26.426 0.080 / 0.053 0.812 / 0.901
De-smokeGCN (Chen et al., 2020) 6.558 / 5.118 24.340 / 26.125 0.067 / 0.059 0.725 / 0.837
MSBDN (Dong et al., 2020) 4.861 / 3.999 25.532 / 27.496 0.058 / 0.051 0.834 / 0.905
SVPNet (Wang et al., 2024a) 4.509 / 3.743 26.663 / 28.121 0.052 / 0.047 0.892 / 0.915
CGAN-DC (Chen et al., 2019) 5.746 / 4.599 25.429 / 26.969 0.059 / 0.054 0.782 / 0.832
MARS-GAN (Hong et al., 2023) 4.863 / 4.592 26.096 / 27.276 0.055 / 0.051 0.888 / 0.910
SSIM-PAN (Sidorov et al., 2020) 5.382 / 4.282 25.092 / 27.003 0.061 / 0.052 0.799 / 0.877
RSTN (Wang et al., 2023) 4.321 / 3.532 26.361 / 28.395 0.053 / 0.047 0.854 / 0.915
LGUTransformer (Wang et al., 2024b) 16.862 / 5.538 17.866 / 27.532 0.130 / 0.075 0.189 / 0.895
Table 6: Training on Cholec80 (synthetic smoke) and evaluation on VASST-desmoke (real smoke).
AODNet (Li et al., 2017) 9.609 / 7.774 18.621 / 21.526 0.132 / 0.109 0.729 / 0.798
GCANet (Chen et al., 2019) 9.803 / 7.042 19.338 / 22.220 0.119 / 0.097 0.747 / 0.801
De-smokeGCN (Chen et al., 2020) 8.370 / 7.585 20.654 / 22.088 0.109 / 0.102 0.716 / 0.788
MSBDN (Dong et al., 2020) 8.277 / 7.327 20.360 / 22.153 0.114 / 0.103 0.769 / 0.810
SVPNet (Wang et al., 2024a) 8.105 / 7.797 21.196 / 21.712 0.110 / 0.109 0.806 / 0.793
CGAN-DC (Chen et al., 2019) 8.177 / 7.525 20.663 / 21.988 0.113 / 0.105 0.770 / 0.805
MARS-GAN (Hong et al., 2023) 7.808 / 7.444 21.444 / 22.110 0.107 / 0.104 0.792 / 0.811
SSIM-PAN (Sidorov et al., 2020) 8.350 / 7.600 20.338 / 21.805 0.115 / 0.106 0.761 / 0.806
RSTN (Wang et al., 2023) 7.963 / 7.533 21.034 / 22.065 0.111 / 0.107 0.763 / 0.813
LGUTransformer (Wang et al., 2024b) 7.357 / 7.352 22.108 / 22.111 0.105 / 0.104 0.812 / 0.818
Table 7: Training using synthetic smoke and evaluation using real smoke on Cholec80 (non-overlap videos).
@@
AODNet (Li et al., 2017) 12.551 / 14.814 0.440 / 0.484 10.449 / 9.434
GCANet (Chen et al., 2019) 7.795 / 11.566 0.407 / 0.403 9.866 / 9.331
De-smokeGCN (Chen et al., 2020) 4.857 / 5.575 0.384 / 0.376 10.885 / 9.569
MSBDN (Dong et al., 2020) 11.838 / 12.176 0.421 / 0.411 10.299 / 9.026
SVPNet (Wang et al., 2024a) 12.812 / 12.134 0.416 / 0.402 10.579 / 9.387
CGAN-DC (Chen et al., 2019) 10.258 / 9.945 0.413 / 0.404 9.517 / 9.039
MARS-GAN (Hong et al., 2023) 13.999 / 11.680 0.416 / 0.413 13.336 / 13.884
SSIM-PAN (Sidorov et al., 2020) 10.539 / 11.093 0.411 / 0.403 9.811 / 11.999
RSTN (Wang et al., 2023) 13.868 / 12.652 0.420 / 0.414 12.245 / 16.910
LGUTransformer (Wang et al., 2024b) 9.126 / 11.236 0.417 / 0.416 13.734 / 12.471
Table 8: Training on VASST-desmoke (real smoke) and evaluation on Hamlyn Dataset (real smoke).
@@
AODNet (Li et al., 2017) 36.549 / 30.515 0.331 / 0.322 12.418 / 11.851
GCANet (Chen et al., 2019) 13.354 / 26.555 0.324 / 0.294 11.512 / 10.983
De-smokeGCN (Chen et al., 2020) 33.236 / 28.007 0.329 / 0.325 29.448 / 18.802
MSBDN (Dong et al., 2020) 26.992 / 25.198 0.384 / 0.322 14.224 / 10.944
SVPNet (Wang et al., 2024a) 39.862 / 30.325 0.401 / 0.317 16.245 / 13.306
CGAN-DC (Chen et al., 2019) 14.052 / 16.699 0.352 / 0.336 16.616 / 13.216
MARS-GAN (Hong et al., 2023) 33.722 / 30.601 0.377 / 0.369 15.011 / 16.381
SSIM-PAN (Sidorov et al., 2020) 11.936 / 16.047 0.357 / 0.298 16.613 / 20.297
RSTN (Wang et al., 2023) 35.629 / 30.232 0.412 / 0.330 16.442 / 12.645
LGUTransformer (Wang et al., 2024b) 35.260 / 29.706 0.295 / 0.294 15.300 / 12.135
Table 9: Training on Cholec80 (synthetic smoke) and evaluation on Hamlyn Dataset (real smoke).
@@
AODNet (Li et al., 2017) 30.292 / 31.168 0.426 / 0.432 11.074 / 11.834
GCANet (Chen et al., 2019) 19.366 / 26.102 0.385 / 0.293 11.628 / 11.083
De-smokeGCN (Chen et al., 2020) 13.920 / 17.647 0.330 / 0.315 11.628 / 11.467
MSBDN (Dong et al., 2020) 26.630 / 27.559 0.409 / 0.332 12.152 / 11.882
SVPNet (Wang et al., 2024a) 29.065 / 28.375 0.406 / 0.267 14.358 / 12.292
CGAN-DC (Chen et al., 2019) 25.175 / 24.127 0.408 / 0.315 10.445 / 10.741
MARS-GAN (Hong et al., 2023) 29.811 / 27.287 0.409 / 0.311 16.443 / 15.136
SSIM-PAN (Sidorov et al., 2020) 24.605 / 26.001 0.395 / 0.312 10.852 / 19.337
RSTN (Wang et al., 2023) 26.305 / 28.347 0.414 / 0.328 13.875 / 14.217
LGUTransformer (Wang et al., 2024b) 23.974 / 26.497 0.383 / 0.346 16.786 / 17.448
@@ 4.3 Comparison of Quantitative Results
The quantitative results on VASST-desmoke and VASST-desmoke-C are presented in Tables 4, 5 and 6. They demonstrate that incorporating SurgiATM effectively improves model performance, regardless of the backbone architecture or loss function. However, for SVPNet in Table 6, SurgiATM yields no improvement in SSIM. We argue that, compared to other desmoking methods, SVPNet (Wang et al., 2024a) employs an energy function (Wang et al., 2018) for restoration. This is a physics-guided approach similar to ours, and its three physics-related parameters are predicted by three independent U-Nets. As a result, incorporating an extra physics-guided component (our SurgiATM) into SVPNet may slightly disrupt the internal physical consistency, leading to the observed deterioration in SSIM.
For Cholec80, we report the evaluation results in Table 7. Our SurgiATM effectively improves most model performances in terms of the fog-related metric (i.e., FADE score). In terms of the external evaluation on the Hamlyn Dataset, as shown in Tables 9 and 8, for the defogging metric FADE, employing our SurgiATM enhances most methods, except for the traditional dehazing method AODNet in Table 9. We attribute this to two factors: 1) the simple architecture and insufficient trainable weights of AODNet (i.e., only five convolutional layers) compared to other methods, and 2) the non-Laplacian error distribution of AODNet when trained on Cholec80 with synthetic smoke, as reported in Table 2. In contrast, although GCANet and SSIM-PAN also exhibit non-Laplacian error distributions in Cholec80, and CGAN-DC and MARS-GAN in VASST-desmoke, their performances are effectively enhanced owing to both their well-designed architectures and the strong generalizability of SurgiATM.
On the other hand, SurgiATM demonstrates limited effectiveness in terms of BRISQUE and NIQE across the ten existing methods in Tables 7, 8 and 9. The main reason is that these two metrics are designed to measure similarity to natural images, which may not generalize well to surgical endoscopic frames due to the substantial domain gap between them (Wang et al., 2024b). The scarcity of domain relevent non-referenced IQA metrics is exactly an underexplored research problem in the area of surgical smoke removal. However, addressing this limitation is beyond the scope of the current study, and developing reliable, domain-specific metrics for surgical videos remains a promising direction for future research.
@@ 4.4 Comparison of Qualitative Results
In this section, we present qualitative comparisons across several representative baselines: De-smokeGCN, GCANet, MARS-GAN, and SVPNet, each selected for different reasons. Specifically, De-smokeGCN (Chen et al., 2020) is a classic DNN-based desmoking method introduced in 2020; GCANet (Chen et al., 2019) is designed for natural dehazing and deraining; MARS-GAN represents an adversarial-learning-based approach; and SVPNet (Wang et al., 2024a), published in 2024, uses a physics-based loss function tailored for surgical desmoking. We train the models on Cholec80 with synthetic smoke and evaluate them on VASST-desmoke, Cholec80, and the Hamlyn Dataset with real surgical smoke, challenging their stability and generalizability.
As shown in the error maps in Fig. 7, SurgiATM effectively reduces desmoking errors in a computationally efficient and interpretable manner, making it well-suited for deployment on edge devices and for stable surgical desmoking. For the natural dehazing method GCANet, it causes color distortion in Fig. 7(a), which is corrected to a certain degree by using SurgiATM. In addition, the proposed approach can mitigate the grid artifact (see the first two rows in Fig. 7), demonstrating promising stability. As shown in Fig. 8, the desmoking results of the baselines are improved to varying degrees. For instance, De-smokeGCN and GCANet can produce more stable results with SurgiATM by reducing the grid artifacts (see Fig. 8(a)-(d) of De-smokeGCN) and correcting the restoration colors (see Fig. 8(d) of GCANet). Moreover, SurgiATM helps baselines eliminate more smoke, as shown in Fig. 8(a) for MARS-GAN and Fig. 8(c) for SVPNet, GCANet, and MARS-GAN.
5 Discussion